A process is running with a mean of 100 microns and standard deviation of 5
microns; specification limits are ±5. What percent falls above 105 microns?

Answer by ewatrrr(24785)

mean of 100 microns and standard deviation of 5
z = (x -100)/5
z = (105-100)/5 = 5/5 = 1
p(x > 5) = p(z > 1) = .1587 0r 15.87%